# DPCore: Dynamic Prompt Coreset for Continual Test-Time Adaptation

### v3

@@ Abstract
Continual Test-Time Adaptation (CTTA) seeks to adapt source pre-trained models to continually changing, unseen target domains. While existing CTTA methods assume structured domain changes with uniform durations, real-world environments often exhibit dynamic patterns where domains recur with varying frequencies and durations. Current approaches, which adapt the same parameters across different domains, struggle in such dynamic conditions—they face convergence issues with brief domain exposures, risk forgetting previously learned knowledge, or misapplying it to irrelevant domains. To remedy this, we propose DPCore, a method designed for robust performance across diverse domain change patterns while ensuring computational efficiency. DPCore integrates three key components: Visual Prompt Adaptation for efficient domain alignment, a Prompt Coreset for knowledge preservation, and a Dynamic Update mechanism that intelligently adjusts existing prompts for similar domains while creating new ones for substantially different domains. Extensive experiments on four benchmarks demonstrate that DPCore consistently outperforms various CTTA methods, achieving state-of-the-art performance in both structured and dynamic settings while reducing trainable parameters by 99% and computation time by 64% compared to previous approaches. Code is available at: https://github.com/zybeich/DPCore.
@@ 2 Preliminaries and Problem Formulation
In this section, we introduce Vision Transformers (ViTs) and formally define the CTTA problem across both CSC and our proposed CDC settings. Due to page limits, detailed discussions on related work are provided in Appendix A.
@@ 2.3 Continual Dynamic Change: A new CTTA setup
The CTTA setting in previous works (Wang et al., 2022a; Liu et al., 2023b, 2024) typically assumes that target domains change in a structured manner: each domain has uniform length and changes occur at regular intervals (Fig. 1 Top). However, real-world scenarios present more dynamic patterns where the same domain may appear multiple times with varying durations, (Fig. 1 Bottom). To systematically study this setting, we simulate dynamic environments using a Dirichlet distribution with parameter $\delta$. As demonstrated in Fig. 4a, smaller values of $\delta$ result in domain transitions more closely resembling the conventional CSC setting. As $\delta$ increases, domain changes become more frequent and unpredictable. For our main experiments, we set $\delta=1$ to keep a balance between structured and completely random domain changes. Various $\delta$ and other distributions are discussed in Appendix E.
@@ 3.1 Visual Prompt Adaptation (VPA)
$d(\Gamma^{S},\Gamma^{T}_{t}(\bm{p}))=\|\bm{\mu}^{S}-\bm{\mu}^{T}_{t}(\bm{p})\|_{2}+\|\bm{\sigma}^{S}-\bm{\sigma}^{T}_{t}(\bm{p})\|_{2}$ (4)
@@
$\bm{p}^{*}=\mathop{\arg\min}\limits_{\bm{p}}\;d(\Gamma^{S},\Gamma_{t}(\bm{p}))$ (5)
@@ 3.3 Dynamic Update to the Prompt Coreset
$\bm{p}_{w}$ $:=\sum_{j=1}^{K}w^{j}\bm{p}^{j},$ (6)
$\mathrm{where}\;\;w^{j}$ $=\frac{\exp(-d(\Gamma_{t},\Gamma^{j})/\tau)}{\sum_{l=1}^{K}\exp(-d(\Gamma_{t},\Gamma^{l})/\tau)}$
@@
$\bm{p}^{j}\leftarrow\bm{p}^{j}+\alpha w^{j}(\bm{p}_{t}-\bm{p}^{j}),\;\Gamma^{j}\leftarrow\Gamma^{j}+\alpha w^{j}(\Gamma_{t}-\Gamma^{j}).$ (7)
@@
3.4 Analysis
We further provide theoretical insight into why DPCore maintains consistent performance across domain changes. Inspired by online K-Means (Duda & Hart, 1973), we analyze a simplified version of DPCore where batches naturally cluster into distinct groups based on their distances. Under the assumption that these clusters are well-separated, we prove three key propositions: 1) (Prop C.1) DPCore correctly assigns batches to their respective clusters regardless of the sequence length, 2) (Prop C.2) these assignments remain correct regardless of the order in which batches arrive, and 3) (Prop C.3) the learned prompts are independent of batch order. Together, these properties demonstrate that DPCore can effectively manage domain knowledge even when domains appear in arbitrary orders with varying frequencies—a crucial advantage in dynamic environments. Detailed analysis is provided in Appendix C.4.
@@ 4.1 Experimental Setup
Methods Compared. We compare our method against strong CTTA baselines including Tent (Wang et al., 2021), SAR (Niu et al., 2023), CoTTA (Wang et al., 2022a), VDP (Gan et al., 2023), DePT (Gao et al., 2023), RoTTA (Yuan et al., 2023), EcoTTA (Song et al., 2023), ViDA (Liu et al., 2023b), BDG (Yang et al., 2024), and C-MAE (Liu et al., 2024). Following their protocols, we conduct several warm-up iterations on labeled source data (e.g., ImageNet) for VDP, DePT, ViDA, and EcoTTA to ensure optimal performance. Implementation details are in Appendix B.
CSC and CDC Settings. We evaluate DPCore across the CSC and CDC settings. The CSC setting follows the implementation by (Wang et al., 2022a). In the experiment under CDC setting, we employ Dirichlet Distribution with $\delta=1$ to generate the data stream from all domains dynamically. Other distributions are discussed in Appendix E.
Implementation Details. To ensure consistency and comparability in our experiments, we follow the implementation details specified in CoTTA (Wang et al., 2022a) and ViDA (Liu et al., 2023b). For classification tasks, we use the ViT-Base model (Wightman, 2019), while for segmentation tasks, we employ the pre-trained Segformer-B5 model (Xie et al., 2021). Following the protocol established in (Zhang et al., 2022), we determine hyperparameters using the four disjoint validation corruptions provided with ImageNet-C and CIFAR10-C (Hendrycks & Dietterich, 2019). We utilize the AdamW optimizer (Loshchilov & Hutter, 2017; Kingma & Ba, 2014) with a learning rate of 0.01, and select the number of prompt tokens at $L=8$ and the threshold at $\rho=0.8$. The source statistics are computed using 300 examples from the corresponding source domain. For other datasets that lack validation sets, we apply these same hyperparameters without additional tuning to align with practical testing conditions, where hyperparameters are selected before accessing any target data. More details of DPCore’s implementation are available in Appendix C.3. We ensure a fair comparison by maintaining the same hyperparameters across both the CSC and CDC settings for all methods.
@@ 4.2 Main Results
We present experimental results on ImageNet-to-ImageNet-C and Cityscapes-to-ACDC across both CSC and CDC settings. Results for CIFAR10/100-to-CIFAR10/100 are provided in Appendix D.3 due to space limit.
@@
To further validate DPCore’s robustness, we conducted additional experiments using 10 random domain orders, following the evaluation protocol of CoTTA (Wang et al., 2022a). DPCore demonstrates consistent performance, achieving an average error rate of 40.2% with an average coreset size of 13.9 elements in Fig. 3c. Details are available in Sec. D.2.
@@
Table 2: Average mIoU score (%) for Cityscapes-to-ACDC across both CSC and CDC settings. The same target domains are repeated three rounds. Full results in Appendix D.4.
@@
[…]
Cityscapes-to-ACDC in CSC and CDC. Beyond classification, we evaluate DPCore on real-world semantic segmentation using the Cityscapes-to-ACDC dataset, where performance is measured by mean Intersection over Union (mIoU, %). As shown in Table 2, in the CSC setting following (Liu et al., 2023b, 2024) where domain sequences repeat three times, DPCore shows consistent improvement in mIoU across cycles (62.1 $\rightarrow$ 62.6 $\rightarrow$ 62.8), outperforming previous SOTA ViDA by 0.6%. In the more challenging CDC setting, while Tent and EcoTTA perform worse than the source model and ViDA’s improvement drops from +5.2% to +2.5%, DPCore maintains robust performance (from +5.8% to +4.3%) while adapting its coreset size from 5 elements in CSC to 7 in CDC to handle increased domain complexity. Fine-grained results are available in Appendix D.4.
@@ 4.3 Ablation Studies
We provide comprehensive analyses and ablation studies for DPCore on ImageNet-to-ImageNet-C in the CSC setting. Additional results are available in Appendix F.
@@
Effect of test batch size. To comprehensively assess the impact of test batch size, we evaluate various CTTA methods across batch sizes ranging from 1 to 256. The results illustrated in Fig. 5d reveal a consistent pattern across methods: regardless of their objective functions (entropy, consistency, or distribution alignment), they maintain stable performance with sufficiently large batch sizes but exhibit increased error rates as batch sizes decrease. In the extreme case of single-sample adaptation (batch size one), all methods demonstrate significantly degraded performance, falling below the source model’s baseline. This observation aligns with findings reported in recent literature (Yuan et al., 2023; Song et al., 2023). Based on these empirical findings and consistent with prior work, we standardize the batch size to 64 for fair comparison across methods. Moreover, DPCore maintains exceptional stability for batch sizes above 16 and shows minimal degradation for smaller sizes (8 and 4) while outperforming other methods. For extremely small batch sizes (e.g. 1), we introduce DPCore-B, which uses a buffer zone to accumulate samples before updates. With a buffer size of 64, DPCore-B achieves a 41.2% error rate even with single-sample batches, demonstrating its practical utility (details in Appendix F.1).
Effect of number of source examples. We evaluate DPCore’s sensitivity to source data volume by varying the number of source examples used for computing statistics from 0 to 10k. As shown in Fig. 5c, DPCore maintains effective adaptation even with 50 examples. For all main experiments, we randomly select 300 unlabeled source examples, notably different from methods like VDP, DePT, EcoTTA, and ViDA that require the entire labeled source dataset for parameter initialization. We further explore an extreme scenario where source data is completely inaccessible (shown at source data = 0 in Fig. 5c). For ImageNet-C experiments, using 300 unlabeled STL10 (Coates et al., 2011) images as reference still achieves an error rate of 45.6%, improving over the source model by 10.2%. Details are provided in Appendix F.2.
@@
ViDA 7.13M 11 1 2.8 43.4
[…]
DPCore’s computation and memory efficiency. We analyze computational complexity across methods in Table 4, comparing learnable parameters, forward/backward propagation counts, and relative computation time (normalized to Tent). DPCore achieves efficiency by introducing only 0.08M parameters (0.1% of model parameters), whereas ViDA requires 7.13M parameters (89× more) and 55.6% more computation time. While this analysis primarily focuses on test-time computation, DPCore also maintains efficiency in pre-adaptation: unlike previous methods which require extensive source data warm-up (e.g., EcoTTA needs 3 epochs of ImageNet training for parameter initialization), DPCore requires minimal preparation, involving only forwarding 300 source examples and computing their statistics. This shows DPCore’s efficiency in both the preparation and adaptation phases. Detailed analysis is in Appendix F.3.

### v4

@@ Abstract
Continual Test-Time Adaptation (CTTA) seeks to adapt source pre-trained models to continually changing, unseen target domains. While existing CTTA methods assume structured domain changes with uniform durations, real-world environments often exhibit dynamic patterns where domains recur with varying frequencies and durations. Current approaches, which adapt the same parameters across different domains, struggle in such dynamic conditions—they face convergence issues with brief domain exposures, risk forgetting previously learned knowledge, or misapplying it to irrelevant domains. To remedy this, we propose DPCore, a method designed for robust performance across diverse domain change patterns while ensuring computational efficiency. DPCore integrates three key components: Visual Prompt Adaptation for efficient domain alignment, a Prompt Coreset for knowledge preservation, and a Dynamic Update mechanism that intelligently adjusts existing prompts for similar domains while creating new ones for substantially different domains. Extensive experiments on four benchmarks demonstrate that DPCore consistently outperforms various CTTA methods, achieving state-of-the-art performance in both structured and dynamic settings while reducing trainable parameters by 99% and computation time by 64% compared to previous approaches. Code is available at https://github.com/yunbeizhang/DPCore.
@@ 1 Introduction
2 Related Works
Test-Time Adaptation (TTA) (Wang et al., 2021; Iwasawa & Matsuo, 2021; Liang et al., 2020; Gao et al., 2023; Schneider et al., 2020; Mirza et al., 2022, 2023; Niu et al., 2024; Liang et al., 2023; Sun et al., 2020; Xiao & Snoek, 2024; Xiao et al., 2023; Mehra et al., 2024b) enhances pre-trained model performance using unlabeled data at test time, without access to the original training phase. TTA techniques fall into two main categories based on their use of source data. The first adjusts models through self-supervised losses like entropy minimization (Wang et al., 2021; Niu et al., 2023) and consistency maximization (Wang et al., 2022a; Liu et al., 2023b). The second involves preliminary steps using source data: either by extracting source characteristics such as statistics or features (Mirza et al., 2023; Niu et al., 2024; Zhang et al., 2024)) or by warming up injected parameters on source data before adaptation (Lee et al., 2024; Song et al., 2023; Liu et al., 2023b; Gao et al., 2023).
Crucially, both approaches are source-free adaptations since they operate without source data during adaptation. While our method falls into the second category, it significantly reduces source dependency compared to existing methods: DPCore requires only 300 unlabeled source examples to compute statistics, in contrast to DePT (Gao et al., 2023), ViDA (Liu et al., 2023b), VDP (Gan et al., 2023) and BeCoTTA (Lee et al., 2024) which need the entire source dataset for warm-up. Moreover, as demonstrated in Fig. 5c, our approach maintains effectiveness even when source data is completely unavailable.
Continual Test-Time Adaptation (CTTA) (Wang et al., 2022a; Niu et al., 2022; Niloy et al., 2024; Liu et al., 2023a; Boudiaf et al., 2022; Yuan et al., 2023; Hoang et al., 2024; Press et al., 2023) tackles the challenge of continually changing target domains. This paradigm was first introduced by CoTTA (Wang et al., 2022a), which we refer to as Continual Structured Change (CSC), where domains change in a structured manner with clear boundaries and uniform lengths. While several subsequent works (Niu et al., 2023; Boudiaf et al., 2022; Gong et al., 2022; Yuan et al., 2023; Lee et al., 2024) explore TTA challenges such as mixed domains, label imbalance, temporal correlation, and gradual shifts, they still primarily focus on structured domain progressions or static domain scenarios. However, real-world environments often exhibit more dynamic patterns where domains recur with varying frequencies and durations—a critical aspect that existing setups often overlook. To address this gap, we introduce the Continual Dynamic Change (CDC) setting, where domains can appear multiple times with irregular durations, better reflecting realistic deployment scenarios.
The dynamic nature of CDC environments exposes fundamental limitations in current CTTA approaches (see Fig. 4b). Most existing methods employ a single-model strategy, updating the same set of parameters across all encountered domains (Wang et al., 2022a; Liu et al., 2023b; Hoang et al., 2024; Niu et al., 2022). While this approach works reasonably well in structured settings, it struggles significantly in dynamic environments due to three critical issues: convergence difficulties with brief domain exposures, catastrophic forgetting of previously learned knowledge, and negative transfer between dissimilar domains. Although some alternatives exist—such as PeTTA (Hoang et al., 2024), which uses memory components, or RDumb (Press et al., 2023), which employs periodic resets—these still fundamentally rely on single-model adaptation. In contrast, our DPCore addresses these limitations through a multi-modeling strategy via a dynamic prompt coreset. This approach learns and manages distinct prompts for different domain groups, enabling DPCore to preserve and reuse knowledge for similar domains while isolating learning for novel ones, thus providing robust adaptation across the varied domain changes characteristic of CDC settings.
@@ 3 Preliminaries and Problem Formulation
In this section, we introduce Vision Transformers (ViTs) and formally define the CTTA problem across both CSC and our proposed CDC settings.
@@ 3.3 Continual Dynamic Change: A new CTTA setup
The CTTA setting in previous works (Wang et al., 2022a; Liu et al., 2023b, 2024) typically assumes that target domains change in a structured manner: each domain has uniform length and changes occur at regular intervals (Fig. 1 Top). However, real-world scenarios present more dynamic patterns where the same domain may appear multiple times with varying durations, (Fig. 1 Bottom). To systematically study this setting, we simulate dynamic environments using a Dirichlet distribution with parameter $\delta$. As demonstrated in Fig. 4a, smaller values of $\delta$ result in domain transitions more closely resembling the conventional CSC setting. As $\delta$ increases, domain changes become more frequent and unpredictable. For our main experiments, we set $\delta=1$ to keep a balance between structured and completely random domain changes. Various $\delta$ and other distributions are discussed in Appendix D.
@@ 4.1 Visual Prompt Adaptation (VPA)
$d(\Gamma^{S},\Gamma^{T}_{t}(\bm{p}))=\|\bm{\mu}^{S}-\bm{\mu}^{T}_{t}(\bm{p})\|_{2}+\|\bm{\sigma}^{S}-\bm{\sigma}^{T}_{t}(\bm{p})\|_{2}\,.$ (4)
@@
$\bm{p}^{*}=\mathop{\arg\min}\limits_{\bm{p}}\;d(\Gamma^{S},\Gamma_{t}(\bm{p}))\,,$ (5)
@@ 4.3 Dynamic Update to the Prompt Coreset
[…]
@@
$\bm{p}^{j}\leftarrow\bm{p}^{j}+\alpha w^{j}(\bm{p}_{t}-\bm{p}^{j}),\;\Gamma^{j}\leftarrow\Gamma^{j}+\alpha w^{j}(\Gamma_{t}-\Gamma^{j}).$
@@
4.4 Theoretical Analysis
We further provide theoretical insight into why DPCore maintains consistent performance across domain changes. Inspired by online K-Means (Duda & Hart, 1973), we analyze a simplified version of DPCore where batches naturally cluster into distinct groups based on their distances. Under the assumption that these clusters are well-separated, we prove three key propositions: 1) (Prop B.1) DPCore correctly assigns batches to their respective clusters regardless of the sequence length, 2) (Prop B.2) these assignments remain correct regardless of the order in which batches arrive, and 3) (Prop B.3) the learned prompts are independent of batch order. Together, these properties demonstrate that DPCore can effectively manage domain knowledge even when domains appear in arbitrary orders with varying frequencies—a crucial advantage in dynamic environments. Detailed analysis is provided in Appendix B.4.
@@ 5.1 Experimental Setup
Methods Compared. We compare our method against strong CTTA baselines including Tent (Wang et al., 2021), SAR (Niu et al., 2023), CoTTA (Wang et al., 2022a), VDP (Gan et al., 2023), DePT (Gao et al., 2023), RoTTA (Yuan et al., 2023), EcoTTA (Song et al., 2023), ViDA (Liu et al., 2023b), BDG (Yang et al., 2024), and C-MAE (Liu et al., 2024). Following their protocols, we conduct several warm-up iterations on labeled source data (e.g., ImageNet) for VDP, DePT, ViDA, and EcoTTA to ensure optimal performance. Implementation details are in Appendix A.
CSC and CDC Settings. We evaluate DPCore across the CSC and CDC settings. The CSC setting follows the implementation by (Wang et al., 2022a). In the experiment under CDC setting, we employ Dirichlet Distribution with $\delta=1$ to generate the data stream from all domains dynamically. Other distributions are discussed in Appendix D.
Implementation Details. To ensure consistency and comparability in our experiments, we follow the implementation details specified in CoTTA (Wang et al., 2022a) and ViDA (Liu et al., 2023b). For classification tasks, we use the ViT-Base model (Wightman, 2019), while for segmentation tasks, we employ the pre-trained Segformer-B5 model (Xie et al., 2021). Following the protocol established in (Zhang et al., 2022), we determine hyperparameters using the four disjoint validation corruptions provided with ImageNet-C and CIFAR10-C (Hendrycks & Dietterich, 2019). We utilize the AdamW optimizer (Loshchilov & Hutter, 2017; Kingma & Ba, 2014) with a learning rate of 0.01, and select the number of prompt tokens at $L=8$ and the threshold at $\rho=0.8$. The source statistics are computed using 300 examples from the corresponding source domain. For other datasets that lack validation sets, we apply these same hyperparameters without additional tuning to align with practical testing conditions, where hyperparameters are selected before accessing any target data. More details of DPCore’s implementation are available in Appendix B.3. We ensure a fair comparison by maintaining the same hyperparameters across both the CSC and CDC settings for all methods.
@@ 5.2 Main Results
We present experimental results on ImageNet-to-ImageNet-C and Cityscapes-to-ACDC across both CSC and CDC settings. Results for CIFAR10/100-to-CIFAR10/100 are provided in Appendix C.3 due to space limit.
@@
To further validate DPCore’s robustness, we conducted additional experiments using 10 random domain orders, following the evaluation protocol of CoTTA (Wang et al., 2022a). DPCore demonstrates consistent performance, achieving an average error rate of 40.2% with an average coreset size of 13.9 elements in Fig. 3c. Details are available in Sec. C.2.
@@
Table 2: Average mIoU score (%) for Cityscapes-to-ACDC across both CSC and CDC settings. The same target domains are repeated three rounds. Full results in Appendix C.4.
@@
DPCore 61.4 60.8 60.7 61.0 +4.3
Cityscapes-to-ACDC in CSC and CDC. Beyond classification, we evaluate DPCore on real-world semantic segmentation using the Cityscapes-to-ACDC dataset, where performance is measured by mean Intersection over Union (mIoU, %). As shown in Table 2, in the CSC setting following (Liu et al., 2023b, 2024) where domain sequences repeat three times, DPCore shows consistent improvement in mIoU across cycles (62.1 $\rightarrow$ 62.6 $\rightarrow$ 62.8), outperforming previous SOTA ViDA by 0.6%. In the more challenging CDC setting, while Tent and EcoTTA perform worse than the source model and ViDA’s improvement drops from +5.2% to +2.5%, DPCore maintains robust performance (from +5.8% to +4.3%) while adapting its coreset size from 5 elements in CSC to 7 in CDC to handle increased domain complexity. Fine-grained results are available in Appendix C.4.
@@ 5.3 Ablation Studies
We provide comprehensive analyses and ablation studies for DPCore on ImageNet-to-ImageNet-C in the CSC setting. Additional results are available in Appendix E.
@@
Effect of test batch size. To comprehensively assess the impact of test batch size, we evaluate various CTTA methods across batch sizes ranging from 1 to 256. The results illustrated in Fig. 5d reveal a consistent pattern across methods: regardless of their objective functions (entropy, consistency, or distribution alignment), they maintain stable performance with sufficiently large batch sizes but exhibit increased error rates as batch sizes decrease. In the extreme case of single-sample adaptation (batch size one), all methods demonstrate significantly degraded performance, falling below the source model’s baseline. This observation aligns with findings reported in recent literature (Yuan et al., 2023; Song et al., 2023). Based on these empirical findings and consistent with prior work, we standardize the batch size to 64 for fair comparison across methods. Moreover, DPCore maintains exceptional stability for batch sizes above 16 and shows minimal degradation for smaller sizes (8 and 4) while outperforming other methods. For extremely small batch sizes (e.g. 1), we introduce DPCore-B, which uses a buffer zone to accumulate samples before updates. With a buffer size of 64, DPCore-B achieves a 41.2% error rate even with single-sample batches, demonstrating its practical utility (details in Appendix E.1).
Effect of number of source examples. We evaluate DPCore’s sensitivity to source data volume by varying the number of source examples used for computing statistics from 0 to 10k. As shown in Fig. 5c, DPCore maintains effective adaptation even with 50 examples. For all main experiments, we randomly select 300 unlabeled source examples, notably different from methods like VDP, DePT, EcoTTA, and ViDA that require the entire labeled source dataset for parameter initialization. We further explore an extreme scenario where source data is completely inaccessible (shown at source data = 0 in Fig. 5c). For ImageNet-C experiments, using 300 unlabeled STL10 (Coates et al., 2011) images as reference still achieves an error rate of 45.6%, improving over the source model by 10.2%. Details are provided in Appendix E.2.
@@
ViDA 93.70M 11 1 2.8 43.4
Ours 0.08M 3.1 1.1 1.8 39.9
DPCore’s computation and memory efficiency. We analyze computational complexity across methods in Table 4, comparing learnable parameters, forward/backward propagation counts, and relative computation time (normalized to Tent). DPCore achieves remarkable efficiency by introducing only 0.08M parameters (0.1% of model parameters), whereas ViDA requires 93.70M parameters (7.13M additional parameters, 89× more) and 55.6% more computation time. While this analysis primarily focuses on test-time computation, DPCore also maintains efficiency in pre-adaptation: unlike previous methods which require extensive source data warm-up (e.g., EcoTTA needs 3 epochs of ImageNet training for parameter initialization), DPCore requires minimal preparation, involving only forwarding 300 source examples and computing their statistics. This shows DPCore’s efficiency in both the preparation and adaptation phases. Detailed analysis is in Appendix E.3.
Fixed-Size Coreset Management. We explore fixed-size coreset strategies for scenarios with strict memory constraints. We evaluate two approaches to maintain a fixed-size coreset on ImageNet-C with K=15: once the number of prompts exceeds K, we either 1) discard the oldest prompt or 2) merge the most similar prompts by averaging those with the smallest statistics distance. As shown in Table 5, both strategies significantly improve upon the source model, with merging performing slightly better than discarding. Since the number of domains is typically unknown in real-world scenarios and the coreset grows only when encountering truly unseen domains, our default approach allows K to evolve naturally.
Table 5: Memory management strategies for fixed-size coreset (K=15) on ImageNet-to-ImageNet-C across 10 different orders.
Algo. Source DPCore
(flexible K) (K=15, discard) (K=15, merge)
Err. Mean $\downarrow$ 55.8 40.2 42.3 41.1